OpenStudy (superhelp101):

Simplify. Express the product as a radical expression. open parentheses x to the fourth power close parentheses to the 3 fifths power

OpenStudy (superhelp101):

@texaschic101

OpenStudy (superhelp101):

@FriedRice @ganeshie8

OpenStudy (superhelp101):

help me anyone!

OpenStudy (superhelp101):

@iPwnBunnies

OpenStudy (superhelp101):

@ganeshie8 plzz

OpenStudy (superhelp101):

@jim_thompson5910

ganeshie8 (ganeshie8):

\[\Large \left(x^4\right)^{\frac{3}{5}}\]

ganeshie8 (ganeshie8):

like that ?

OpenStudy (superhelp101):

yes

OpenStudy (superhelp101):

x fifth root of x squared x to the fourth power fifth root of x cubed x cubed fifth root of x squared x squared fifth root of x squared these are my options

ganeshie8 (ganeshie8):

familiar with any exponent properties ?

OpenStudy (superhelp101):

yes

ganeshie8 (ganeshie8):

\[\huge \left(x^4\right)^{\frac{3}{5}} = x^{\frac{12}{5}} = x^{\frac{5*2+2}{5}}\]

ganeshie8 (ganeshie8):

\[\huge = x^{\frac{5*2}{5}}. x^{\frac{2}{5}}\]

ganeshie8 (ganeshie8):

\[\huge = x^{2}. \sqrt[5]{x^2}\]

OpenStudy (superhelp101):

Awwwwww, thank you so much! You're very sweet. Thank you! :)

ganeshie8 (ganeshie8):

yw!